# 2INFN, Laboratori Nazionali del Gran Sasso, I-67100 Assergi, Italy 3Universität Heidelberg, Zentrum für Astronomie (ZAH), Institut für Theoretische Astrophysik, Albert Ueberle Str. 2, 69120, Heidelberg, Germany 4Universität Heidelberg, Interdisziplinäres Zentrum für Wissenschaftliches Rechnen, Heidelberg, Germany

### v1

@@ 3.1.2 Delay times and formation channels
Figure 4 shows the delay time (i.e., the time between the formation of the stellar binary system and its merger) for BHNS mergers in our fiducial Pop. III and Pop. II models. We compare these distributions with the $\propto t^{-1}$ trend from Dominik et al. (2012). This is the expected scaling of $t_{\rm del}$ if we consider orbital periods with a nearly log-uniform distribution (as in kro1 and log1, see Table 1). Deviations from this trend suggest a final semi-major axis distribution significantly different compared to the initial one (Peters 1964; Sana et al. 2012). In general, we find a good agreement between our simulations and the $\propto t^{-1}$ trend.
We point out a deviation from the trend for Pop. III BHNS mergers with $t_{\rm del}>10^{3}\,\rm Myr$, which is due to a majority of systems with semi-major axes larger than $\sim 10\,\rm R_{\odot}$. In Table 4, we report the fraction of BHNS mergers happening through the formation channels previously described in Sec. 2.6, and summarized in Table 2. We find that Pop. III BHNSs predominantly form via channel 2, which involves only stable mass transfer episodes and, as a result, does not significantly reduce the semi-major axis.
By contrast, Pop. II BHNS mergers show an excess at short delay times ($t_{\rm del}\sim 15\,\rm Myr$). This is due to the dominance of Pop. II BHNSs forming via channel 1, which shrinks semi-major axes effectively via common envelope.
Figure 4: Distribution of delay times of BHNS mergers for the fiducial models of Pop. III (continuous blue line) and Pop. II stars (dotted blue line). The gray dashed line shows the $\propto t^{-1}$ trend from Dominik et al. (2012).
@@ 3.1.3 Primary mass and mass ratio
Figure 5 shows the primary mass (i.e., the BH mass) distribution of BHNS mergers. Pop. II BHNS mergers show a peak at $m_{1}\lesssim 15\,{\rm M}_{\odot}$ (Giacobbo & Mapelli 2018; Broekgaarden et al. 2021). After this peak, the number of BHNS mergers decreases with $m_{1}$ up to $\sim 60\,{\rm M}_{\odot}$. BHNS mergers from Pop. III stars, instead, have a flatter distribution in $m_{1}$ between 5 and $\sim 47\,{\rm M}_{\odot}$ and show a peak around $22\,{\rm M}_{\odot}$. The different trends of these distributions are related to the dependence of the mass-radius relation on the metallicity $Z$ of the stars. In fact, regardless of the formation channel of BHNS mergers, Pop. II star binaries with $m_{1,\rm ZAMS}>40\,{\rm M}_{\odot}$ will tend to prematurely merge unless $a_{\rm ZAMS}\gtrsim 10^{3}\,\rm R_{\odot}$. Pop. III star binaries with high $m_{1,\rm ZAMS}$ will instead expand less during their evolution, and avoid premature mergers if $a_{\rm ZAMS}\gtrsim 400\,\rm R_{\odot}$. As reported in Table 4, the lack of a prominent peak at $m_{1}<20\,{\rm M}_{\odot}$ and the choice of a top-heavier IMF, lead to a lower merger efficiency for Pop. III BHNSs $\eta_{\rm BHNS}$.
Figure 6 shows the dependence of the primary mass of BHNS mergers on the delay time $t_{\rm del}$. In the case of Pop. III stars, high-mass systems merge preferentially at large delay times. This means that, at low-redshift, we might be able to observe GW signals from BHNSs with $m_{1}>20\,{\rm M}_{\odot}$ deriving from the evolution of metal-free binaries. In fact, more than $85\%$ of Pop. III BHNS mergers have $m_{1}>20\,{\rm M}_{\odot}$, while the fraction reduces to $\sim 11\%$ for Pop. II. As a consequence, even though at low redshifts the MRD of Pop. II BHNSs is about one order of magnitude larger than the one of Pop. III BHNSs (see Fig. 3), we have eight times more mergers from Pop. III BHNSs with $m_{1}>20\,{\rm M}_{\odot}$ compared to Pop. II. The contribution of Pop. III BHNS mergers to the currently observed GW events will be further discussed in Sec. 4.2.
Similarly to Pop. III, Pop. II BHNSs with massive primary BHs tend to merge with longer delay times (bottom panel of Figure 6). In fact, Pop. II binaries with large $m_{\rm ZAMS,1}$ are generally able to avoid premature mergers only when their initial semi-major axis is wide enough, resulting in longer delay times.
@@
Figure 6: Distribution of primary mass of BHNS mergers in three ranges of $t_{\rm del}$. We show the results for the fiducial models of Pop. III (top), and Pop. II progenitors (bottom). We represent in violet mergers with $t_{\rm del}<20\,\rm Myr$, in blue-green mergers with $20<t_{\rm del}<10^{3}\,\rm Myr$, and in green mergers with $t_{\rm del}>10^{3}\,\rm Myr$.
Finally, Fig. 7 shows the distributions of mass ratio $q$ of the simulated BHNS mergers. Pop. III stars yield a distribution that peaks at $q<0.1$ and extends up to $q\sim 0.3$; in fact, as we have seen, metal-free stars favor BHNS mergers with a large primary mass, resulting in small values of $q$. In contrast, Pop. II BHNSs merge with a mass ratio that peaks at $q\sim 0.2$ and extends to larger values, up to $q\sim 0.35$. In accordance with both Broekgaarden et al. (2021) and Giacobbo & Mapelli (2018), this is a consequence of the peak at $m_{1}<15\,{\rm M}_{\odot}$ (Fig. 5) and of the lower mass gap, between $2$ and $5\,{\rm M}_{\odot}$. These distinguishing features are particularly interesting because they can as well contribute to a distinction between GW events from Pop. III and Pop. II binaries (see Sec. 4.2).
Figure 7: Distribution of mass ratios $q$ of BHNS mergers for the fiducial models of Pop. III (continuous blue line) and Pop. II (dotted blue line).
@@ 3.2.1 Merger rate density
Figure 8 shows the MRD of BNSs from Pop. III and Pop. II stars. For Pop. III stars, we take into account all the simulated models reported in Table 1. In the case of Pop. II stars, instead, we only consider the fiducial model kro1. As we did for BHNSs, we compare these MRDs with the one derived by Iorio et al. (2023), which includes higher metallicity stars.
The MRD of BNSs from Pop. III stars peaks at $z\sim 15$, for all considered models. Here, the different models we simulated span five orders of magnitude $\left[\mathcal{R}_{\rm BNS}(z\sim 15)\sim 10^{-3}-15\,\rm Gpc^{-3}\,yr^{-1}\right]$. Because of the extremely short BNS delay times (see Sec. 3.2.2 and Fig. 9), the MRDs from Pop. III stars peak in the vicinity of the peak of the chosen star formation rate density. The trend of all Pop. III MRDs at $z<5$ is also due to the convolution of the small delay times ($<50\,\rm Myr$) with the chosen star formation rate density $\psi(z^{\prime},Z)$ (see Fig. 2). Model kro1 produces the highest MRD for Pop. III stars, with $\mathcal{R}_{\rm BNS}(z)\sim 15\,\rm Gpc^{-3}\,yr^{-1}$ at $z\sim 15$. The peak of our fiducial Pop. III model (log1) is an order of magnitude lower. In fact, kro1 produces less massive BNS progenitors involved in premature mergers. As discussed in Sec. 3.1.1, models adopting the initial orbital period distribution from Stacy & Bromm (2013) also result in lower MRDs in this case.
The BNS MRD from Pop. II stars dominates over the one from Pop. III stars up to $z\sim 10$, where it reaches a value of $\mathcal{R}_{\rm BNS}(z\sim 10)\sim 5\,\rm Gpc^{-3}\,yr^{-1}$. There, the MRD of Pop. III BNSs from model kro1 is within the same order of magnitude. As pointed out for Pop. III BNSs, also Pop. II BNS mergers have extremely short delay times, leading to a drop in the MRD at $z<5$.
Figure 1 shows that Pop. III stars with $m_{\rm ZAMS}\lesssim 20\,{\rm M}_{\odot}$ are more compact with respect to Pop. II stars. As a consequence, they produce, on the one hand, less premature mergers and, on the other, a higher BNS merger efficiency ($\eta_{\rm BNS}\sim 10^{-4}\,{\rm M}_{\odot}^{-1}$ for model kro1). However, as shown in Table 5, convolving the merger efficiency with the star formation rate density models yields a MRD from Pop. II BNSs that exceeds that of Pop. III by at least one order of magnitude at $z<10$.
For comparison, we show that the MRD of BNSs born from Pop. II-I progenitors (Iorio et al. 2023) peaks at $z\sim 2$ ($\mathcal{R}_{\rm BNS}(z\sim 2)\sim 10^{3}\,\rm Gpc^{-3}\,yr^{-1}$), consistently with the peak of the star formation rate density from Madau & Fragos (2017). The contribution of BNSs born from more metal-rich progenitors dominates the MRDs below $z\sim 11$.
@@
Table 5: Same as Table 3 but for the merger rate density of BNSs from Pop. III and Pop. II stars.
@@ 3.2.2 Delay times and formation channels
Figure 9 shows the distribution of delay times of BNS mergers for the fiducial models of Pop. III and Pop. II stars. Both populations result in distributions deviating significantly from the $\propto t^{-1}$ trend predicted by Dominik et al. (2012). The delay times of BNS mergers show a peak at $\sim 20\,\rm Myr$, and a common trend up to $t_{\rm del}\sim 10^{3}\,\rm Myr$, after which we observe an excess of BNS mergers from Pop. II stars.
The peak at $\sim 20\,\rm Myr$ is the result of very efficient orbital shrinking after one (or more) common envelope event(s). As a matter of fact, Table 6 shows that BNS mergers take place when a binary system undergoes at least one common envelope during its evolution (channels 1,3,4,5). As a consequence, more than $98\,\%$ of the merging BNSs have $a<1\,\rm R_{\odot}$.
The excess of Pop. II systems at $t_{\rm del}>10^{3}\,\rm Myr$ can be also explained by looking at Table 6. In fact, half of the BNS mergers from Pop. II stars form through channel 3 (see Table 2), which favors binaries with $a_{\rm ZAMS}\sim 200-5\times 10^{4}\,\rm R_{\odot}$. Indeed, Fig. 1 shows that Pop. II stars with $m_{\rm ZAMS}\sim 8-20\,{\rm M}_{\odot}$ tend to expand to larger radii with respect to Pop. III stars in the same mass range. Consequently, only binary stars with large initial separations will avoid to merge prematurely. We find that a fraction of the binaries evolving through channel 3 yields BNSs with $a>2\,\rm R_{\odot}$, that will need $\geq 10^{3}\,\rm Myr$ to merge.
Lastly, Fig. 9 points out the lower merger efficiency of Pop. II BNSs, with an evident lack of systems merging with $50<t_{\rm del}<10^{3}\,\rm Myr$. In Table 6 we show that, for the chosen fiducial models, the merger efficiency of Pop. II BNSs is $\sim 1.4$ times lower than the one of Pop. III BNSs.
@@ 4 Discussion
In the following, we discuss the impact of the initial conditions on the properties of BHNS mergers. We then compute the detectability-conditioned Bayes factors for three GW events comparing a Pop. III to a Pop. II origin.
@@ 4.1 Impact of initial conditions on distributions of $m_{1}$ and $q$
Figure 10 shows the distribution of the primary mass $m_{1}$ of BHNS mergers for all the simulated models of Pop. III stars (Table 1). As a comparison, we also show the distributions for the two Pop. II star models assuming a Kroupa IMF (kro1, kro5; Kroupa 2001).
@@
All Pop. III models produce BHNS mergers with a primary mass between $\sim 6\,{\rm M}_{\odot}$ and $\sim 50\,{\rm M}_{\odot}$, with the exception of log3, which pushes $m_{1}$ above the upper mass gap, up to $\sim 400\,{\rm M}_{\odot}$. In fact, the initial mass-ratio distribution adopted in model log3 (”sorted”; see Table 1) favors the formation of binaries with $m_{1,\rm ZAMS}\gtrsim 200\,{\rm M}_{\odot}$ and $m_{2,\rm ZAMS}\lesssim 25\,{\rm M}_{\odot}$. These high-mass binaries typically have initial separations $a_{\rm ZAMS}>10^{3}\,\rm R_{\odot}$ and merge predominantly through channel 3 (Table 2), which favors unequal-mass systems (Costa et al. 2023).
@@
Models with initial orbital parameters from Stacy & Bromm (2013) (log2, log5, kro5, lar5, top5) tend to produce BHNS mergers with $m_{1}\gtrsim 15\,{\rm M}_{\odot}$ and peaking at $m_{1}\gtrsim 20\,{\rm M}_{\odot}$. These models yield less BHNS mergers because of the wide initial semi-major axes; merging BHNSs have $a_{\rm ZAMS}\sim 10^{3}-10^{5}\,\rm R_{\odot}$ and evolve mainly through channel 3. Differently from log3, these models rely on an initial mass-ratio distribution from Stacy & Bromm (2013) with $q_{\rm min}=0.1$, which favors the formation of BHNSs with large $m_{\rm 1,ZAMS}$, but does not allow systems with $m_{\rm 1,ZAMS}>200\,{\rm M}_{\odot}$.
@@
Finally, models based on Sana et al. (2012) yield a flatter $m_{1}$ distribution, with a sharp peak around $m_{1}\sim 22\,{\rm M}_{\odot}$. We point out that model kro1 produces a factor three more mergers with respect to the Pop. III fiducial model log1.
@@
Consequently, all the simulated Pop. III star models show a nearly flat distribution of BH mass for $m_{1}>20\,{\rm M}_{\odot}$. Moreover, in Sec. 3.1.3, we have seen that these massive systems usually merge with large delay times. Thus, it is important to take into account the contribution of Pop. III massive BHNS mergers also at low redshifts (see Sec. 4.2).
Pop. II BHNS mergers from model kro1 show a peak at $m_{1}\sim 10\,{\rm M}_{\odot}$ and follow a decreasing trend up to $\sim 60\,{\rm M}_{\odot}$ (see Sec. 3.1.3).
@@
As in the case of Pop. III, Pop. II BHNS mergers from model kro5 (Stacy & Bromm 2013) mostly have $m_{1}>15\,{\rm M}_{\odot}$. Model kro5 is thus able to reproduce massive BHNS mergers without the help of stellar dynamics (Rastello et al. 2020; Arca Sedda 2020; Arca sedda et al. 2024). Nevertheless, because of their larger radial expansion with respect with Pop. III stars, Pop. II binaries are also able to produce more BHNS mergers with low-mass primaries.
@@
Figure 11 shows the mass-ratio $q$ distributions for all the simulated models. Starting from Pop. III stars, we notice that models with initial orbital periods from Sana et al. (2012) peak at $q\sim 0.1$ and extend up to $q\sim 0.3$. Models based on Stacy & Bromm (2013), instead, get to a maximum of $q\sim 0.2$. We point out again model log3, producing BHNS mergers with extremely low mass ratios; this feature was already found and discussed by Costa et al. (2023) in the case of binary BHs. Moving on to Pop. II BHNS mergers, we see that model kro1 produces a peak at $q\sim 0.2$ and extends to larger mass ratios with respect to its metal-free counterpart (see also Sec. 3.1.3). Due to its primary mass distribution, model kro5 shows intermediate features between Pop. III BHNS mergers with initial parameter distributions from Stacy & Bromm (2013) and from Sana et al. (2012).

### v2

@@ 3.1.2 Delay times and formation channels
Figure 4 shows the delay time (i.e., the time between the formation of the stellar binary system and its merger) for BHNS mergers in our fiducial Pop. III and Pop. II models. We compare these distributions with the $\propto t^{-1}$ trend from dominik2012. This is the expected scaling of $t_{\rm del}$ if we consider orbital periods with a nearly log-uniform distribution (as in kro1 and log1, see Table 1). Deviations from this trend suggest a final semimajor axis distribution that is significantly different from the initial one (peters1964; sana2012). In general, we find a good agreement between our simulations and the $\propto t^{-1}$ trend.
We point out a deviation from the trend for Pop. III BHNS mergers with $t_{\rm del}>10^{3}\,\rm Myr$, which is due to a majority of systems with semimajor axes larger than $\sim 10\,\rm R_{\odot}$. In Table 4, we report the fraction of BHNS mergers happening through the formation channels previously described in Sec. 2.6, and summarized in Table 2. We find that Pop. III BHNSs predominantly form via channel 2, which involves only stable mass transfer episodes and, as a result, does not significantly reduce the semimajor axis. The second most common pathway is channel 3, which is particularly relevant for systems with wide initial separations.
By contrast, Pop. II BHNS mergers show an excess at short delay times ($t_{\rm del}\sim 15\,\rm Myr$). This is due to the large fraction ($\sim{45}\%)$ of Pop. II BHNSs forming via channel 1, which shrinks semimajor axes effectively via the common envelope evolution, whereas channel 2 and 3 tend to have longer timescales.
Figure 4: Distribution of delay times of BHNS mergers for the fiducial models of Pop. III (continuous blue line) and Pop. II stars (dotted blue line). The dashed gray line shows the $\propto t^{-1}$ trend from dominik2012.
@@ 3.1.3 Primary mass and mass ratio
Figure 5 shows the primary mass (i.e., the BH mass) distribution of BHNS mergers. Pop. II BHNS mergers show a peak at $m_{1}\lesssim 15\,{\rm M}_{\odot}$ (giacobbo2018; floor2021). After this peak, the number of BHNS mergers decreases with $m_{1}$ up to $\sim 60\,{\rm M}_{\odot}$. The BHNS mergers from Pop. III stars, instead, have a flatter distribution in $m_{1}$ between 5 and $\sim 47\,{\rm M}_{\odot}$ and show a peak around $22\,{\rm M}_{\odot}$. The different trends of these distributions are related to the dependence of the mass-radius relation on the metallicity, $Z$, of the stars. In fact, regardless of the formation channel of BHNS mergers, Pop. II star binaries with $m_{1,\rm ZAMS}>40\,{\rm M}_{\odot}$ will tend to prematurely merge unless $a_{\rm ZAMS}\gtrsim 10^{3}\,\rm R_{\odot}$. Population III star binaries with high $m_{1,\rm ZAMS}$ will instead expand less during their evolution, and avoid premature mergers if $a_{\rm ZAMS}\gtrsim 400\,\rm R_{\odot}$. As is reported in Table 4, the lack of a prominent peak at $m_{1}\leq 20\,{\rm M}_{\odot}$ and the choice of a top-heavier IMF lead to a lower merger efficiency for Pop. III BHNSs, $\eta_{\rm BHNS}$.
Figure 6 shows the cumulative distribution of delay times for two primary mass bins: light ($m_{1}\leq 20\,{\rm M}_{\odot}$) and heavy ($m_{1}>20\,{\rm M}_{\odot}$) primary BHs. High-mass Pop. III BHNS mergers are associated with longer delay times ($t_{\rm del}>10^{3}\,\rm Myr$) compared to light systems. This means that, at low redshifts, we might be able to observe GW signals from BHNSs with $m_{1}>20\,{\rm M}_{\odot}$ deriving from the evolution of metal-free binaries. In fact, more than $85\%$ of Pop. III BHNS mergers have $m_{1}>20\,{\rm M}_{\odot}$, while the fraction reduces to $\sim 11\%$ for Pop. II. As a consequence, even though at low redshifts the MRD of Pop. II BHNSs is about one order of magnitude larger than the one of Pop. III BHNSs (see Fig. 3), we have eight times more mergers from Pop. III BHNSs with $m_{1}>20\,{\rm M}_{\odot}$ compared to Pop. II. The contribution of Pop. III BHNS mergers to the currently observed GW events is discussed further in Sec. 4.2.
The smaller population of Pop. II BHNS mergers with $m_{1}>20\,{\rm M}_{\odot}$ also tends to occur at long delay times, following a trend similar to Pop. III (see bottom panel of Fig. 6). In both cases, in fact, binaries with massive $m_{\rm ZAMS,1}$ avoid premature mergers only when their initial separations are wide, leading to longer $t_{\rm del}$.
The delay times of light BHNS mergers, instead, have different trends for the two stellar populations. The subdominant Pop. III BHNS mergers with $m_{1}\leq 20\,{\rm M}_{\odot}$ have smaller delay times with respect to Pop. II BHNSs, owing to the greater compactness of Pop. III progenitors at low $m_{\rm 1,ZAMS}$ (see Fig. 1), which facilitates the formation of close compact binaries.
@@
Figure 6: Cumulative distribution of the delay times, $t_{\rm del}$, of BHNS mergers in two ranges of primary mass, $m_{1}$. We represent in violet mergers with $m_{1}\leq 20\,{\rm M}_{\odot}$ and in blue-green mergers with $m_{1}>20\,{\rm M}_{\odot}$. We show the results for the fiducial models of Pop. III (left), and Pop. II progenitors (right).
Finally, Fig. 7 shows the distributions of the mass ratio, $q$, of the simulated BHNS mergers. Pop. III stars yield a distribution that peaks at $q<0.1$ and extends up to $q\sim 0.3$; in fact, as we have seen, metal-free stars favor BHNS mergers with a large primary mass, resulting in small values of $q$. In contrast, Pop. II BHNSs merge with a mass ratio that peaks at $q\sim 0.2$ and extends to larger values, up to $q\sim 0.35$. In accordance with both floor2021 and giacobbo2018, this is a consequence of the peak at $m_{1}<15\,{\rm M}_{\odot}$ (Fig. 5) and of the lower mass gap, between $2$ and $5\,{\rm M}_{\odot}$. These distinguishing features are particularly interesting because they can also contribute to a distinction between GW events from Pop. III and Pop. II binaries (see Sec. 4.2).
Figure 7: Distribution of mass ratios, $q$, of BHNS mergers for the fiducial models of Pop. III (continuous blue line) and Pop. II (dotted blue line).
@@ 3.2.1 Merger rate density
Figure 8 shows the MRD of BNSs from Pop. III and Pop. II stars. For Pop. III stars, we take into account all the simulated models reported in Table 1. In the case of Pop. II stars, instead, we only consider the fiducial model kro1. As we did for BHNSs, we compare these MRDs with the one derived by sevn2023, which includes higher-metallicity stars.
The MRD of BNSs from Pop. III stars peaks at $z\sim 15$, for all considered models. Here, the different models we simulated span five orders of magnitude: $\left[\mathcal{R}_{\rm BNS}(z\sim 15)\sim 10^{-3}-15\,\rm Gpc^{-3}\,yr^{-1}\right]$. Because of the extremely short BNS delay times (see Sec. 3.2.2 and Fig. 9), the MRDs from Pop. III stars peak in the vicinity of the peak of the chosen star formation rate density. At $z<5$, their behavior is driven by the convolution of the short delay times ($50\,\mathrm{Myr}$) with the chosen $\psi(z^{\prime},Z)$ (see Fig. 2). Since Pop. III BNSs lack systems with long delay times, their MRDs rapidly decline once the underlying star formation rate density drops, without extending significantly beyond the star formation rate density cutoff.
Model kro1 produces the highest MRD for Pop. III stars, with $\mathcal{R}_{\rm BNS}(z)\sim 15\,\rm Gpc^{-3}\,yr^{-1}$ at $z\sim 15$. The peak of our fiducial Pop. III model (log1) is an order of magnitude lower. In fact, kro1 produces less massive BNS progenitors involved in premature mergers. As is discussed in Sec. 3.1.1, models adopting the initial orbital period distribution from sb2013 also result in lower MRDs in this case. In Appendix C, we take into account the impact of Pop. III star formation rate density models on the MRDs.
The BNS MRD from Pop. II stars dominates over the one from Pop. III stars up to $z\sim 10$, where it reaches a value of $\mathcal{R}_{\rm BNS}(z\sim 10)\sim 5\,\rm Gpc^{-3}\,yr^{-1}$. There, the MRD of Pop. III BNSs from model kro1 is within the same order of magnitude. As was pointed out for Pop. III BNSs, Pop. II BNS mergers also have extremely short delay times, leading to a drop in the MRD at $z<5$.
Figure 1 shows that Pop. III stars with $m_{\rm ZAMS}\lesssim 20\,{\rm M}_{\odot}$ are more compact with respect to Pop. II stars. As a consequence, they produce, on the one hand, less premature mergers and, on the other, a higher BNS merger efficiency ($\eta_{\rm BNS}\sim 10^{-4}\,{\rm M}_{\odot}^{-1}$ for model kro1). However, as is shown in Table 5, convolving the merger efficiency with the star formation rate density models yields a MRD from Pop. II BNSs that exceeds that of Pop. III by at least one order of magnitude at $z<10$.
For comparison, we show that the MRD of BNSs born from Pop. II-I progenitors (sevn2023) peaks at $z\sim 2$ ($\mathcal{R}_{\rm BNS}(z\sim 2)\sim 10^{3}\,\rm Gpc^{-3}\,yr^{-1}$), consistently with the peak of the star formation rate density from madau2017. The contribution of BNSs born from more metal-rich progenitors dominates the MRDs below $z\sim 11$.
@@
Table 5: Merger rate density of BNSs from Pop. III and Pop. II stars. [^5] Same as Table 3 but for the MRD of BNSs from Pop. III and Pop. II stars.
@@
Table 6: Formation channels of BNS mergers. [^6] Same as Table 4 but for BNS mergers.
@@ 3.2.2 Delay times and formation channels
Figure 9 shows the distribution of delay times of BNS mergers for the fiducial models of Pop. III and Pop. II stars. Both populations result in distributions deviating significantly from the $\propto t^{-1}$ trend predicted by dominik2012. The delay times of BNS mergers show a peak at $\sim 20\,\rm Myr$, and a common trend up to $t_{\rm del}\sim 10^{3}\,\rm Myr$, after which we observe an excess of BNS mergers from Pop. II stars.
The peak at $\sim 20\,\rm Myr$ is the result of very efficient orbital shrinking after one (or more) common envelope event(s). As a matter of fact, Table 6 shows that BNS mergers take place when a binary system undergoes at least one common envelope event during its evolution (channels 1, 3, 4, and 5). As a consequence, more than $98\,\%$ of the merging BNSs have $a<1\,\rm R_{\odot}$ at the time of formation of the second neutron star.
The excess of Pop. II systems with $t_{\rm del}>10^{3}\,\rm Myr$ can also be explained by looking at Table 6. In fact, half of the BNS mergers from Pop. II stars form through channel 3 (see Table 2), which favors binaries with $a_{\rm ZAMS}\sim 200-5\times 10^{4}\,\rm R_{\odot}$. Indeed, Fig. 1 shows that Pop. II stars with $m_{\rm ZAMS}\sim 8-20\,{\rm M}_{\odot}$ tend to expand to larger radii with respect to Pop. III stars in the same mass range. Consequently, only binary stars with large initial separations will avoid to merge prematurely. We find that a fraction of the binaries evolving through channel 3 yields BNSs with $a>2\,\rm R_{\odot}$, which will need $\geq 10^{3}\,\rm Myr$ to merge.
Lastly, Fig. 9 points out the lower merger efficiency of Pop. II BNSs, with an evident lack of systems merging with $50<t_{\rm del}<10^{3}\,\rm Myr$. In Table 6 shows that, for the chosen fiducial models, the merger efficiency of Pop. II BNSs is $\sim 1.4$ times lower than the one of Pop. III BNSs.
@@ 4 Discussion
In the following, we discuss the impact of the initial conditions on the properties of BHNS mergers. We then compute the detectability-conditioned Bayes factors for three GW events comparing a Pop. III origin to a Pop. II one.
@@ 4.1 Impact of initial conditions on distributions of $m_{1}$ and $q$
Figure 10 shows the distribution of the primary mass, $m_{1}$, of BHNS mergers for all the simulated models of Pop. III stars (Table 1). As a comparison, we also show the distributions for the two Pop. II star models assuming a Kroupa IMF (kro1, kro5; kroupa2001).
@@
All Pop. III star models produce BHNS mergers with a primary mass of between $\sim 6\,{\rm M}_{\odot}$ and $\sim 50\,{\rm M}_{\odot}$, with the exception of log3, which pushes $m_{1}$ above the upper mass gap, up to $\sim 400\,{\rm M}_{\odot}$. In fact, the initial mass-ratio distribution adopted in model log3 (“sorted;” see Table 1) favors the formation of binaries with $m_{1,\rm ZAMS}\gtrsim 200\,{\rm M}_{\odot}$ and $m_{2,\rm ZAMS}\lesssim 25\,{\rm M}_{\odot}$. These high-mass binaries typically have initial separations of $a_{\rm ZAMS}>10^{3}\,\rm R_{\odot}$ and merge predominantly through channel 3 (Table 2), which favors unequal-mass systems (costa2023).
@@
Models with initial orbital parameters from sb2013 (log2, log5, kro5, lar5, and top5) tend to produce BHNS mergers that have $m_{1}\gtrsim 15\,{\rm M}_{\odot}$ and peak at $m_{1}\gtrsim 20\,{\rm M}_{\odot}$. These models yield fewer BHNS mergers because of the wide initial semimajor axes; merging BHNSs have $a_{\rm ZAMS}\sim 10^{3}-10^{5}\,\rm R_{\odot}$ and evolve mainly through channel 3. Unlike log3, these models rely on an initial mass-ratio distribution from sb2013 with $q_{\rm min}=0.1$, which favors the formation of BHNSs with large $m_{\rm 1,ZAMS}$, but does not allow systems with $m_{\rm 1,ZAMS}>200\,{\rm M}_{\odot}$.
@@
Finally, models based on sana2012 yield a flatter $m_{1}$ distribution, with a sharp peak around $m_{1}\sim 22\,{\rm M}_{\odot}$. We point out that model kro1 produces more mergers by a factor of three with respect to the Pop. III fiducial model log1.
@@
Consequently, all the simulated Pop. III star models show a nearly flat distribution of BH mass for $m_{1}>20\,{\rm M}_{\odot}$. Moreover, in Sec. 3.1.3, we have seen that these massive systems usually merge with large delay times. Thus, it is important to also take into account the contribution of Pop. III massive BHNS mergers at low redshifts (see Sec. 4.2).
@@
Pop. II BHNS mergers from model kro1 show a peak at $m_{1}\sim 10\,{\rm M}_{\odot}$ and follow a decreasing trend up to $\sim 60\,{\rm M}_{\odot}$ (see Sec. 3.1.3). As in the case of Pop. III stars, Pop. II BHNS mergers from model kro5 (sb2013) mostly have $m_{1}>15\,{\rm M}_{\odot}$. Model kro5 is thus able to reproduce massive BHNS mergers without the help of stellar dynamics (rastello2020; mas2020; mas2024). Nevertheless, because of their larger radial expansion with respect to Pop. III stars, Pop. II binaries are also able to produce more BHNS mergers with low-mass primaries.
@@
Figure 11 shows the mass-ratio $q$ distributions for all the simulated models. Starting from Pop. III stars, we notice that models with initial orbital periods from sana2012 peak at $q\sim 0.1$ and extend up to $q\sim 0.3$. Models based on sb2013, instead, get to a maximum of $q\sim 0.2$. We point out again model log3, producing BHNS mergers with extremely low mass ratios; this feature was already found and discussed by costa2023 in the case of binary BHs. Moving on to Pop. II BHNS mergers, we see that model kro1 produces a peak at $q\sim 0.2$ and extends to larger mass ratios with respect to its metal-free counterpart (see also Sec. 3.1.3). Due to its primary mass distribution, model kro5 shows intermediate features between Pop. III BHNS mergers with initial parameter distributions from sb2013 and from sana2012.